OpenStudy (joannablackwelder):

cos(cos^-1(-3/7)-tan^-1(-12/5)) Any help on finding the exact value of this expression would be greatly appreciated! Thanks!

OpenStudy (freckles):

well first do you remember that cos(a-b)=cos(a)cos(b)+sin(a)sin(b) ?

OpenStudy (joannablackwelder):

Ah, gotcha! I was just thinking about that!

OpenStudy (freckles):

let a=arccos(3/7) and b=arctan(-12/5) so cos(a)=3/7 and tan(b)=-12/5 draw two separate right triangles

OpenStudy (freckles):

let me know if you need anymore help

OpenStudy (joannablackwelder):

Got it, thanks!